Find the cylindrical coordinates of the points where the Cartesian coordinates are :


(√5, 1,2)


1
Expert's answer
2022-01-25T12:12:19-0500

Cylindrical coordinates "(r,\theta,z)."

"r=\sqrt{x^2+y^2}=\sqrt{\sqrt{5}^2+1^2}=\sqrt{6}."

"\theta=arctan{\frac{y}{x}}=arctan{\frac{1}{\sqrt{5}}}=0.4205."

"z=z=2."

Cylindrical coordinates: "(\sqrt{6},0.4205,2)."